Three resistors R1, R2 and R3 are connected in series-parallel with R1 in series with the parallel combination of R2 and R3. The whole combination is connected across a 120 V DC source. Resistors R1, R2 and R3 take 750 W, 250 W and 200 W respectively. Calculate the resistance R2.

1
Expert's answer
2021-12-12T16:40:05-0500

for parallel combination of R2 and R3:

"U_2=U_3"

then:

"U_1+U_2=120" V


"P_2=U_2I_2=250" W

"P_3=U_3I_3=200" W

"I_2\/I_3=250\/200=1.25"


"I_1=I_2+I_3=2.25I_3"


"P_1\/I_1+P_2\/I_2=120"


"750\/(2.25I_3)+250\/(1.25I_3)=120"

"I_3=4.44" A

"I_2=1.25\cdot4.44=5.56" A


"U_2=P_2\/I_2=250\/5.56=44.96" V


"R_2=U_2/I_2=44.96/5.56=8.09\ \Omega"